Circular pipe of 23.9 mm OD and 1.5 mm wall thickness and length 90 mm is used as a fin on a wall at 200°C. The outside is exposed to air at 35°C with a convection coefficient of 15 W/m2K. The conductivity of the material is 45 W/mK. Determine the heat loss from the fin and also the mid section and tip temperature.
Answer #1

iten data For circular fin D: 23.9 mm inside . 23.9 - 2 1 .5 - thickness = 1.5mm AC = (do-d}-] = I [23.9*-20.9?] = 105.5 mm2Tip temp. Ti-Tool To - Too coshme Ti - 35 F 299 - 35 losh (4.61X0.09) TL = 186.75°C]